9. A buffer system is made up of equal concentrations of HAc and Ac─. When H+ ions are added to this buffer, it will resist decrease in pH; in doing so the concentration of HAc will rise. Explain if agree or disagree.




10. Carbonic acid is one of the three buffers used by living beings. Two pka of carbonic acid are 6.35 and 10.33. Write the weak acid and its conjugate based that will make the most effective buffer cellular function. _______________________________ __________________________________

BONUS: Arrange the following in order of increasing energy content: (a) Covalent C-C bond (b) Glucose molecule (c) Protein molecule (d) hydrogen bond
Answer #1

Solution:

9) According to Hendersen-Hasselbalch equation,

pH = pka + log [Ac-] / [HAc]

From above equation, it can be seen that when the concentrations of HAc- and Ac- are equal, then,

pH = pka + log 1

pH = pka

As, [Ac-] / [HAc] goes from 1/10 to 10 the pH changes from pKa + 1 to pKa -1.

Thus, addition of small amount of H+ ion does not affect the pH of buffer. Thus, buffer resists the decrease in pH.

The addition of H+ ion in the buffer causes conversion of Ac- to HAc and hence, the ratio [Ac-]/[HAc] decreases.

H+ + Ac- = HAc

Since, pH of the buffer is calculated as,

pH = pka + log [Ac-] / [HAc]

Thus, on addition of H+, the value of log [Ac-] / [HAc] changes significantly. It resists the decrease in pH on addition of H+ ions.

Therefore, agree with the statement.